OpenStudy (anonymous):

2.5 + 2x = 0.5 solve for x

OpenStudy (anonymous):

2x=0.5-2.5 2x=-2 x=-1

OpenStudy (anonymous):

To answer equations like this you want to get the variable (x in this case) completely alone on one side. I'm using a hypothetical equations, please don't mistake this for your question. But the process is the same. So first you make sure all of the variables are on the same side. If they are not you want to get them there. If you have 4x+5=3−x You want both x to be on the same side. The easiest in this case is to add x to both sides to get rid of the -x on the right. Because in the end it's better to have all positive x. So now you do 4x+5+x=3−x+x Do you see how the -x + x Cancels each other out? So you're left with: 5x+5=3 Second you would do any multiplication or division that effects the whole side x was on. I don't really know how to do good fractions on this site, so I'm going to skip how to do that overall, scroll down to the how to do it to x, hopefully it'll make sense there, it's the same thing as all of this is. Just do the opposite. Then, see if there are any exponentials or exponential roots that effect all of the side x is on. If there are you would do the opposite of what is there (If the whole side is squared you would do square root, if the whole side is rooted you would square.) You do that opposite to both sides. So if you had a new equation of \[(2x+7)^{2}=4\] you would take the square root of both sides to get ride of the root on the side where the x is. (You'd be left with (2x+7)=2. Then you look at any addition or subtraction on the same side as x. In this case the +7. You want to move that. To do so you do the opposite. Since this is addition you would do subtraction to both sides. The equation would look like so 2x+7−7=2−7 Which, when you do the math becomes 2x=−5 Next you would do any multiplication or division that effects x only. In this case it's times 2. So you would divide both sides by 2. Again, I don't know how to do good fractions, so you'll have to bear with me and hopefully it makes sense. Then, finally, as you did before you do exponents or roots that effect only the variable. In short it's: Get the variable on the same side. Whole side exponents / roots Whole side multiplication / division Addition or subtraction Variable multiplication / division Variable exponents / roots.